3. Three Strings A3= 220 HZ, E4=660 HZ, & E5=660 HZ, are sounded simultaneously. a) what are the frequencies of all the beat notes? b) Give the ratios of all the frequencies heard by the ear.
Answer #1

Beats is the difference of two frequencies

Here E4 should be 330 Hz

a) So here beats are= 660-220=440Hz

660-330=330Hz

330-220=110Hz

So total three beats has formed

b) Human ear can hear beats about to 20Hz to 20000Hz

So they all are heard .